OpenStudy (anonymous):

WILL MEDAL AND FAN What is the slope of the equation? y = –5x − 2.

OpenStudy (anonymous):

@Capnkookcx

OpenStudy (igreen):

y = -5x - 2 It's in slope-intercept form. y = mx + b Where 'm' is the slope. So what's the slope?

OpenStudy (anonymous):

@bohotness the equations are different. the 5 in mine is negative

OpenStudy (anonymous):

-5 is the slope?

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

I just did

OpenStudy (anonymous):

can you help again?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

What is the equation of this line of best fit in slope-intercept form? y=-6x+3/4 y=6x+3/4 y=-3/4x+6 y=3/4x+6

OpenStudy (anonymous):

@PizzaLover123

OpenStudy (anonymous):

@iGreen @iGreen @iGreen @iGreen @iGreen @iGreen